# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

### v3

@@ 5.2 Effect of Self-Adaptive Success Rate Sampling
SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by maintaining the randomness of the shaped rewards. To further explore the effect of this mechanism, we use the AntStand task as a case study, analyzing the shaped rewards learned at different stages of training. Figure 4 (bottom) shows the learning curve, while Figure 4 (top) illustrates the distributions of generated rewards over the “height of the ant" feature, one dimension in the state vector.
@@
As learning progresses, the shaped rewards demonstrate two key attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic. Besides, although the sampled rewards fluctuate, their center gradually shows a positive linear correlation with the robot’s height. In early phases, the shaped rewards contain significant randomness due to higher uncertainty. Although these random signals offer limited information, they drive the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies, thus broadens the actions sampled from SAC’s stochastic policy, encouraging exploration and increasing the diversity of collected samples. In later phases, the rewards converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation.
To further investigate the exploration behavior of SASR, we compare the density of the visited states throughout the training in the MountainCar task with four representative exploration strategies: (1) ReLara, perturbating on both rewards and actions; (2) #Explo, rewarding novel states; (3) SAC, entropy-regularized exploration; and (4) NoisyNet (Fortunato et al.,, 2018), perturbing network weights. The density for every 25k steps is shown in Figure 5. We observe that SASR progressively covers a wider range of the state space. From 50k to 100k steps, SASR reaches positions near the goal, driven by the success rate mechanism. In contrast, ReLara covers a similar range to SASR, but is less smooth and takes longer to reach the right side. #Explo shows no clear rightward shift, as it focuses on rewarding novelty without considering the inherent value of states. SAC’s exploration is relatively narrow, making it easier to trap in local optima. NoisyNet shows a narrowing range as network weight perturbations diminish through optimization. Overall, SASR demonstrates more effective exploration and collects valuable samples sooner, leading to faster convergence.
@@ 5.3 Ablation Study
We conduct ablation studies to investigate key components of SASR. We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix D.
@@
(b) Learning performance of SASR with different retention rates $\phi$.
(c) Learning performance of SASR with different bandwidths $h$ of Gaussian kernel.
@@
Sampling from Beta distributions. (Figure 6(a)) We examine a variant of SASR that omits Beta distribution sampling, instead directly using the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$. In the early stages, limited experience makes this success rate an unreliable estimate, and using this fixed, overly confident value can mislead the agent. Furthermore, skipping Beta distribution sampling removes the exploration driven by random rewards, leading to narrower exploration. The results highlight the critical role of Beta distribution sampling for effective learning.
Retention rate $\phi$. (Figure 6(b)) The retention rate controls the estimation of counts, which affects the confidence of Beta distributions. A high retention rate ($\phi=1$) preserves all samples, resulting in a densely populated and redundant state pool, causing the Beta distribution to be prematurely overconfident, which degrades performance. Conversely, a low retention rate ($\phi=0.01$) slows convergence by requiring more iterations to gather sufficient samples. The results suggest that an appropriate retention rate is important.
Bandwidth $h$ of Gaussian kernel. (Figure 6(c)) The bandwidth $h$ controls the smoothness of the kernel functions. Beyond fixed bandwidths, we also test a linearly decreasing configuration ($h:0.5\rightarrow 0.1$), which reflects increasing confidence in KDE. Results indicate that a small bandwidth ($h=0.01$) increases the distance between samples, causing many to have zero estimated density, while a large bandwidth ($h=1$) makes samples indistinguishable due to an overly flat kernel function. Both cases result in suboptimal performance. The decreasing bandwidth setting offers no significant improvement and tends to reduce stability due to inconsistent density estimations.
RFF feature dimensions $M$. (Figure 6(d)) SASR shows relatively low sensitivity to $M$, provided it is large enough to capture the states’ complexity. Results show that values like $M=500,1000,2000$ all yield similar performance levels, while significantly low dimensions, such as $M=50$, degrade performance. This is also suggested in the original RFF study (Rahimi and Recht,, 2007).
Shaped reward weight factor $\lambda$. (Figure 6(e)) Regarding the shaped reward weight factor $\lambda$, we observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$. At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance. At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback and hinder learning. The findings highlight that maintaining a balanced reward scale is important for optimal learning outcomes.
@@ 6 Conclusion and Discussion
In this paper, we propose SASR, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge. SASR achieves an exploration and exploitation balance mechanism by generating shaped rewards from evolving Beta distributions. Experiments demonstrate that this adaptability significantly enhances the agent’s convergence speed. Additionally, the implementation of KDE and RFF formulates a highly efficient, non-parametric, and learning-free approach to deriving Beta distributions. This mechanism also provides a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments. Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability.
While SASR is designed for sparse-reward environments, in settings with dense rewards, the introduction and maintenance of additional shaped rewards could be unnecessary. Exploring the extension of SASR to dense reward scenarios presents a promising direction for further research. Moreover, the derivation of Beta distributions depends heavily on the samples stored in the success and failure buffers. Currently, our method does not account for the relationships or varying importance of different states within the same trajectory, making the algorithm sensitive to the retention rate $\phi$. Therefore, developing an adaptive retention rate or better mechanisms for managing the buffers are crucial avenues for future improvement.
7 Acknowledgement
Appendix A Derivation of Random Fourier Features
We incorporate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to approximate the kernel functions in the KDE process for the SASR algorithm. Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$. RFF approximates the kernel function by projecting the input $k$-dimensional space into a $M$-dimensional feature space using a mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. The RFF-based kernel function is then defined as follows:
@@ Appendix A Derivation of Random Fourier Features
Appendix B Derivation of Computation Complexity
@@
Appendix C Comparison of SASR with and without RFF
@@
Appendix D Supplementary Experimental Results for Ablation Study
We provide the detailed quantitative results of the ablation study in this section.
@@ Appendix D Supplementary Experimental Results for Ablation Study
AntStand $\bm{39.12\pm 2.86}$ $23.25\pm 2.12$
AntFar $\bm{73.92\pm 4.97}$ $37.86\pm 5.16$
HumanStand $\bm{42.63\pm 2.17}$ $9.30\pm 0.083$
HumanKeep $\bm{180.98\pm 4.40}$ $170.51\pm 10.68$
RobotReach $\bm{81.29\pm 6.52}$ $33.76\pm 4.41$
RobotPush $\bm{137.06\pm 12.66}$ $70.37\pm 11.97$
Table 5: Ablation study #2: The average episodic returns and standard errors of SASR with different retention rates.
@@
AntStand $17.98\pm 1.86$ $\bm{39.12\pm 2.86}$ $20.27\pm 2.08$
AntFar $19.95\pm 3.00$ $\bm{73.92\pm 4.97}$ $48.47\pm 5.28$
HumanStand $9.25\pm 0.08$ $\bm{42.63\pm 2.17}$ $19.17\pm 1.52$
HumanKeep $179.10\pm 7.06$ $\bm{180.98\pm 4.40}$ $180.13\pm 6.27$
RobotReach $54.07\pm 8.39$ $\bm{81.29\pm 6.52}$ $35.18\pm 5.95$
RobotPush $3.91\pm 1.22$ $\bm{137.06\pm 12.66}$ $1.88\pm 0.47$
Table 6: Ablation study #3: The average episodic returns and standard errors of SASR with different bandwidths $h$ of Gaussian kernel.
Tasks $h=0.01$ $h=0.1$ $h=0.2$ (default) $h=1$ $h=0.5\rightarrow 0.1$
AntStand $7.94\pm 0.28$ $16.77\pm 1.18$ $\bm{39.12\pm 2.86}$ $12.76\pm 0.40$ $19.00\pm 1.63$
AntFar $11.68\pm 0.96$ $44.31\pm 3.34$ $\bm{73.92\pm 4.97}$ $27.96\pm 2.46$ $72.71\pm 4.57$
HumanStand $9.31\pm 0.06$ $\bm{43.64\pm 1.94}$ $42.63\pm 2.17$ $9.32\pm 0.06$ $13.81\pm 0.93$
HumanKeep $178.51\pm 4.99$ $\bm{181.44\pm 4.73}$ $180.98\pm 4.40$ $178.22\pm 4.97$ $178.04\pm 5.12$
RobotReach $37.03\pm 4.19$ $72.30\pm 5.67$ $\bm{81.29\pm 6.52}$ $35.87\pm 4.57$ $49.47\pm 5.1$
RobotPush $0.40\pm 0.13$ $8.80\pm 4.44$ $\bm{137.06\pm 12.66}$ $0.12\pm 0.06$ $0.34\pm 0.13$
@@
AntStand $3.56\pm 0.12$ $12.08\pm 1.36$ $39.12\pm 2.86$ $\bm{42.52\pm 3.14}$
AntFar $27.76\pm 3.24$ $36.31\pm 2.69$ $\bm{73.92\pm 4.97}$ $61.79\pm 4.63$
HumanStand $9.46\pm 0.05$ $25.42\pm 1.83$ $42.63\pm 2.17$ $\bm{52.70\pm 2.13}$
HumanKeep $178.51\pm 4.84$ $181.81\pm 4.60$ $\bm{180.98\pm 4.40}$ $180.58\pm 4.63$
RobotReach $65.39\pm 6.22$ $38.25\pm 4.37$ $\bm{81.29\pm 6.52}$ $45.98\pm 4.90$
RobotPush $33.58\pm 4.86$ $54.56\pm 6.66$ $\bm{137.06\pm 12.66}$ $45.15\pm 6.15$
@@
AntStand $7.60\pm 0.73$ $17.25\pm 1.58$ $\bm{39.12\pm 2.86}$ $28.43\pm 2.07$ $2.82\pm 0.09$
AntFar $22.51\pm 2.60$ $34.04\pm 3.87$ $\bm{73.92\pm 4.97}$ $67.62\pm 4.89$ $18.05\pm 1.79$
HumanStand $9.53\pm 0.04$ $33.95\pm 1.90$ $\bm{42.63\pm 2.17}$ $30.65\pm 1.65$ $12.23\pm 0.42$
HumanKeep $\bm{182.57\pm 4.32}$ $180.22\pm 4.60$ $180.98\pm 4.40$ $179.74\pm 4.53$ $179.62\pm 4.70$
RobotReach $41.05\pm 4.17$ $\bm{82.33\pm 6.37}$ $81.29\pm 6.52$ $71.30\pm 6.54$ $15.89\pm 2.06$
RobotPush $9.18\pm 1.99$ $118.92\pm 10.79$ $\bm{137.06\pm 12.66}$ $65.51\pm 12.71$ $47.98\pm 9.64$
Appendix E Network Structures and Hyperparameters
E.1 Network Structures

### v4

@@ 5.2 Effect of Self-Adaptive Success Rate Sampling
SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by modulating the randomness of the shaped rewards. To further investigate the effect of this mechanism, we use the AntStand task as a case study, analyzing the shaped rewards learned at different training stages. Figure 4 (bottom) shows the learning curve, while Figure 4 (top) illustrates the distributions of generated rewards over the “height of the ant” feature, a dimension in the state vector.
@@
As learning progresses, the shaped rewards exhibit two key attributes: values transition from random to meaningful, and variance decreases from uncertain to deterministic. Although the sampled rewards fluctuate, their means gradually show a positive linear correlation with the robot’s height. In the early phases, the shaped rewards contain significant randomness due to high uncertainty. While these random signals offer limited information, they drive the agent to take small optimization steps in diverse directions, effectively shifting the policy anchors, expanding the action space sampled from SAC’s stochastic policy, promoting exploration, and increasing sample diversity. In later phases, rewards stabilizes, closely aligning with the robot’s height – a meaningful and intuitive metric – enhancing the agent’s exploitation.
To further investigate SASR’s exploration behavior, we compare the visited state density throughout training in the MountainCar task with five representative exploration strategies: (1) ReLara, which perturbs both rewards and actions; (2) #Explo and (3) RND, which rewards novel states; (4) SAC, which uses entropy-regularized exploration; and (5) NoisyNet (Fortunato et al., 2018), which perturbs network weights. The state density for every 25k steps is shown in Figure 5. We observe that SASR progressively covers a wider range of the state space. From 50k to 100k steps, SASR reaches positions near the goal, driven by its success rate mechanism. In contrast, ReLara and RND cover similar ranges to SASR, but are less smooth and take longer to reach the right side. #Explo shows no clear rightward shift, as it prioritizes novelty and ignores the inherent value of states. SAC’s exploration is relatively narrow, making it prone to getting trapped in local optima. NoisyNet’s range narrows over time as perturbations diminish through optimization. Overall, SASR demonstrates more effective exploration and collects valuable samples sooner, leading to faster convergence.
@@ 5.3 Ablation Study
We conduct ablation studies to investigate key components of SASR. We select six representative tasks and report the experimental results, with quantitative data provided in Appendix A.5.
@@
(b) Learning performance of SASR with shaped reward function $R^{S}(s,a)$ and $R^{S}(s)$.
(c) Learning performance of SASR with different retention rates $\phi$.
@@
Sampling from Beta distributions. (Figure 6(a)) We examine a variant of SASR that omits Beta distribution sampling, instead directly using the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$. In the early stages, limited experience makes this success rate an unreliable estimate, and using a fixed, overly confident value can mislead the agent. Furthermore, skipping Beta distribution sampling eliminates exploration driven by random rewards, leading to narrower exploration. The results highlight the critical role of Beta distribution sampling in effective learning.
Reward function over state-action pair. (Figure 6(b)) We extend SASR with a reward function over state-action pairs, $r(s,a)$. The comparison results show that both settings perform similarly. However, encoding actions into the reward function increases dimensionality, complicating density estimation and correlation assessment. Furthermore, the state and action vectors may have different distributions, potentially reducing KDE estimation accuracy.
Retention rate $\phi$. (Figure 6(c)) The retention rate directly influences the confidence level of Beta distributions. A high retention rate ($\phi=1$) preserves all samples, resulting in a densely populated, redundant state pool, which makes the Beta distribution prematurely overconfident and degrades performance. Conversely, a low retention rate ($\phi=0.01$) slows convergence as more iterations are required to gather sufficient samples. The results suggest that an appropriate retention rate is crucial.
RFF feature dimensions $M$. (Figure 6(d)) SASR shows relatively low sensitivity to $M$, provided it is sufficiently large to capture the complexity of the states. Results show that values like $M=500,1000,2000$ all yield similar performance, while significantly lower dimensions, such as $M=50$, degrade performance. These findings align with the original RFF study (Rahimi & Recht, 2007).
Shaped reward weight factor $\lambda$. (Figure 6(e)) SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$. At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance. At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, potentially overwhelming feedback and hindering learning. The findings emphasize that maintaining a balanced reward scale is important for optimal learning outcomes.
@@ 6 Conclusion and Discussion
In this paper, we propose SASR, a self-adaptive reward shaping algorithm based on success rates to address the sparse-reward challenge. SASR achieves a balance between exploration and exploitation by generating shaped rewards from evolving Beta distributions. Experiments demonstrate that this adaptability significantly improves the agent’s convergence speed. Additionally, the use of KDE and RFF provides a computationally efficient approach to deriving Beta distributions. This mechanism also offers a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments. Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability.
While SASR is designed for sparse-reward environments, in dense-reward settings, the additional shaped rewards may be unnecessary. Extending SASR to such scenarios presents a promising direction for further research. Moreover, the derivation of Beta distributions relies on the samples stored in the success and failure buffers. Currently, our method does not consider the relationships or varying importance of different states within the same trajectory, making it sensitive to the retention rate. Therefore, developing an adaptive retention rate or improved buffer management mechanisms is crucial for future improvement.
Acknowledgement
This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (Award Number: AISG2-RP-2020-016). This research is also partially supported by an Academic Research Grant No. MOE-T2EP20121-0015 from the Ministry of Education in Singapore.
Appendix A Appendix
A.1 Derivation of Random Fourier Features
We incorporate Random Fourier Features (RFF) (Rahimi & Recht, 2007) to approximate the kernel functions in the KDE process for the SASR algorithm. Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$. RFF approximates the kernel function by projecting the input $k$-dimensional space into a $M$-dimensional feature space using a mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. The RFF-based kernel function is then defined as follows:
@@ A.1 Derivation of Random Fourier Features
A.2 Computational Complexity
@@
A.3 Experiments on Time and Space Complexity
A.3.1 Time and Space Complexity Comparison
In this section, we analyze the time and space overhead introduced by SASR and other representative reward-shaping methods. Below, we summarize the computational and memory costs of the RS baselines, introduced by the shaped reward generation.
  • SASR (ours) calculates shaped rewards using RFF, which essentially is matrix operations, without additional networks/models learning processes. Regarding the memory costs, the buffers $D_{S}$ and $D_{F}$ are much smaller than the replay buffer used in the backbone SAC algorithm, due to the retention rate $\phi$. While considering the scalability for larger problems, we have implemented an alternative approach by augmenting the original replay buffer in the backbone SAC algorithm with a success or failure flag. This approach avoids the need for additional buffers.
  • ReLara (Ma et al., 2024a) requires an additional RL agent (of the same scale as the original RL agent) and an additional replay buffer.
  • ROSA (Mguni et al., 2023) involves a competition agent (the same sacle as the original RL agent) and a switching model (a neural network).
  • ExploRS (Devidze et al., 2022) requires learning two parameterized networks: one for a self-supervised reward model and another for the exploration bonus.
  • #Explo (Tang et al., 2017) requires a hash function to discretize the state space and a hash table to store the state-visitation counts.
  • RND (Burda et al., 2018) uses a random network distillation module to compute the intrinsic rewards.
Furthermore, we report the computational and memory costs of SASR and the RS baselines in two tasks: AntStand and Frogger, the results are shown in Table 3 and Table 4, respectively. To provide a more intuitive comparison, we report the relative value normalized to our SASR, in this case, if the value $>1$, it indicates that the baseline is more computationally or memory expensive than SASR, and vice versa.
Table 3: Average maximum memory consumption during the training process, normalized to SASR.
Tasks SASR ReLara ROSA ExploRS #Explo RND
AntStand 1 3.67 4.12 2.05 0.89 0.12
Frogger 1 5.21 4.33 2.64 0.92 0.09
Table 4: Average training time, normalized to SASR.
Tasks SASR ReLara ROSA ExploRS #Explo RND
AntStand 1 1.87 2.12 1.67 1.08 1.11
Frogger 1 1.98 3.17 1.72 1.24 1.06
A.3.2 Comparison of SASR with and without RFF
@@
A.4 Auto Hyperparameter Selection
To improve the robustness and generalization of the SASR algorithm, we propose some potential autonomous hyperparameter selection strategies, mainly designed for the bandwidth $h$ of the kernel function and the RFF feature dimension $M$.
For bandwidth $h$, we can use the empirical formula Silverman’s Rule of Thumb (Wilcox, 2011):
$h=1.06\cdot\sigma\cdot N^{-1/5},$ (26)
or cross-validation to determine the optimal bandwidth.
For the RFF dimension $M$, it is directly related to the bandwidth $h$. After determining $h$, we can use the formula mentioned in the RFF theory to determine $M$:
$M=O(\frac{1}{\epsilon^{2}}\log{\frac{N}{\delta}}),$ (27)
where $\epsilon$ and $\delta$ are the error and confidence parameters. Another method is to compare the Frobenius norm error between the RFF approximated kernel matrix $K^{RFF}$ and the true kernel matrix $K^{Gaussian}$ to select $M$: $\|K^{RFF}-K^{Gaussian}\|_{F}$.
A.5 Supplementary Experimental Results for Ablation Study
In this section, we provide the detailed quantitative results of the ablation study.
@@ A.5 Supplementary Experimental Results for Ablation Study
AntStand 94.92$\pm$0.00 54.48 $\pm$ 1.29
AntFar 139.84$\pm$0.00 92.77 $\pm$ 1.53
HumanStand 79.83$\pm$2.03 9.77 $\pm$ 0.02
HumanKeep 195.77$\pm$0.00 185.00 $\pm$ 0.00
RobotReach 170.18$\pm$0.00 110.29 $\pm$ 2.93
RobotPush 167.14$\pm$0.00 86.82 $\pm$ 0.00
Table 7: Ablation study #2: The average episodic returns and standard errors of SASR with reward function on state-action pair or state only.
Tasks SASR (with $R^{S}(s)$) SASR (with $R^{S}(s,a)$)
AntStand 94.92$\pm$0.00 85.61$\pm$1.30
AntFar 139.84$\pm$0.00 132.49$\pm$2.92
HumanStand 79.83$\pm$2.03 78.93$\pm$0.65
HumanKeep 195.77$\pm$0.00 192.54$\pm$0.16
RobotReach 170.18$\pm$0.00 151.95$\pm$5.74
RobotPush 167.14$\pm$0.00 179.76$\pm$1.66
Table 8: Ablation study #3: The average episodic returns and standard errors of SASR with different retention rates.
@@
AntStand 45.71$\pm$7.57 94.92$\pm$0.00 62.85$\pm$3.49
AntFar 70.07$\pm$3.30 139.84$\pm$0.00 103.65$\pm$2.57
HumanStand 9.88$\pm$0.01 79.83$\pm$2.03 66.46$\pm$2.96
HumanKeep 195.00$\pm$0.00 195.77$\pm$0.00 194.77$\pm$0.10
RobotReach 154.32$\pm$0.89 170.18$\pm$0.00 112.46$\pm$0.90
RobotPush 2.96$\pm$1.97 167.14$\pm$0.00 1.75$\pm$1.24
@@
AntStand 5.21$\pm$0.45 50.68$\pm$6.40 94.92$\pm$0.00 96.80$\pm$8.42
AntFar 98.88$\pm$2.64 72.17$\pm$5.07 139.84$\pm$0.00 129.87$\pm$0.63
HumanStand 9.87$\pm$0.01 78.82$\pm$0.52 79.83$\pm$2.03 77.73$\pm$1.47
HumanKeep 193.84$\pm$0.61 194.71$\pm$0.15 195.77$\pm$0.00 195.86$\pm$0.03
RobotReach 119.09$\pm$26.92 122.89$\pm$4.51 170.18$\pm$0.00 94.87$\pm$15.56
RobotPush 71.67$\pm$27.53 161.70$\pm$11.54 167.14$\pm$0.00 150.20$\pm$8.70
@@
AntStand 35.71$\pm$0.92 59.82$\pm$2.71 94.92$\pm$0.00 75.61$\pm$1.41 3.16$\pm$0.35
AntFar 99.83$\pm$3.25 119.82$\pm$1.18 139.84$\pm$0.00 119.35$\pm$1.80 80.71$\pm$4.74
HumanStand 9.81$\pm$0.02 75.35$\pm$1.06 79.83$\pm$2.03 70.96$\pm$0.51 28.94$\pm$0.46
HumanKeep 194.68$\pm$0.08 194.21$\pm$0.38 195.77$\pm$0.00 193.85$\pm$0.42 194.89$\pm$0.10
RobotReach 131.61$\pm$4.51 154.16$\pm$5.20 170.18$\pm$0.00 169.00$\pm$2.56 74.23$\pm$4.23
RobotPush 13.07$\pm$1.65 193.56$\pm$3.85 167.14$\pm$0.00 178.90$\pm$0.00 192.07$\pm$0.87
Bandwidth $h$ of Gaussian kernel. The bandwidth $h$ controls the smoothness of the kernel functions. Beyond fixed bandwidths, we also test a linearly decreasing configuration ($h:0.5\rightarrow 0.1$), which reflects increasing confidence in KDE. Results indicate that a small bandwidth ($h=0.01$) increases the distance between samples, causing many to have zero estimated density, while a large bandwidth ($h=1$) makes samples indistinguishable due to an overly flat kernel function. Both cases result in suboptimal performance. The decreasing bandwidth setting offers no significant improvement and tends to reduce stability due to inconsistent density estimations.
Table 11: Ablation study #6: The average episodic returns and standard errors of SASR with different bandwidths $h$ of Gaussian kernel.
Tasks $h=0.01$ $h=0.1$ $h=0.2$ (default) $h=1$ $h=0.5\rightarrow 0.1$
AntStand 10.71$\pm$2.52 57.22$\pm$3.87 94.92$\pm$0.00 17.74$\pm$2.53 68.40$\pm$1.54
AntFar 17.58$\pm$2.84 99.80$\pm$4.41 139.84$\pm$0.00 25.83$\pm$8.62 136.49$\pm$4.15
HumanStand 9.89$\pm$0.01 64.47$\pm$1.87 79.83$\pm$2.03 9.90$\pm$0.02 58.79$\pm$2.76
HumanKeep 194.92$\pm$0.02 194.00$\pm$0.57 195.77$\pm$0.00 193.06$\pm$0.46 194.59$\pm$0.18
RobotReach 128.57$\pm$3.83 97.35$\pm$19.12 170.18$\pm$0.00 134.02$\pm$2.02 59.39$\pm$26.02
RobotPush 2.29$\pm$1.62 122.45$\pm$37.58 167.14$\pm$0.00 0.00$\pm$0.00 0.01$\pm$0.01
A.6 Network Structures and Hyperparameters
A.6.1 Network Structures